OpenStudy (anonymous):

Solve for r in this formula S= L - rL where S is the sale price, L is the list price, and r is the discount rate. an item with with a list price of $56.80 goes on sale for $36.92. find the discount rate. an item with a list price of $56.80 goes on sale for $36.92. find the discount rate

OpenStudy (anonymous):

S = L -rL S + rL = L rL = L - S r = (L - S)/L Now substitute the given values and u will get the answer :)